OpenStudy (anonymous):

r(t)= 1)find T(t) at (1,0,0) and explain what this tells you about the motion of particle on the curve at this point 2)find N(t) 3) find an equation of the line tangent to r(t) at this point. 4) find an equation of a plane normal to r(t) at this point. 5) find the curvature at point (1,0,0)

OpenStudy (anonymous):

@TuringTest n @Avva can u guys help me out with this question ?

OpenStudy (anonymous):

sorry Idk how to solve this :(

OpenStudy (turingtest):

this is just a series of formulas...

OpenStudy (anonymous):

@Avva no problem :)

OpenStudy (turingtest):

here is the unit tangent vector formula\[\vec T(t)=\frac{\vec r'(t)}{\|\vec r'(t)\|}\]you should be able to guess what the tangent of a position curve represents

OpenStudy (turingtest):

the unit normal vector has the formula\[\vec N(t)=\frac{\vec T'(t)}{\|\vec T'(t)\|}\]

OpenStudy (anonymous):

for T(t) at (1,0,0) i have \[(0,1/\sqrt{2},0)\] fot T(t) how do i know the motion of a particle ?

OpenStudy (turingtest):

the derivative of position is velocity, so that is the direction of the velocity vector of the particle when it is at (1,0,0)

OpenStudy (anonymous):

T(t) = <-sint, cost,-tant>/\[\sqrt{1+\sec ^{2}}\]

OpenStudy (turingtest):

I have to go, but you are not the only one working on this problem right now why don't you move over to this post and work with amistre64, he can help you :) http://openstudy.com/study#/updates/50ad2c69e4b09749ccabad2f

OpenStudy (anonymous):

@TuringTest : hey thank you so much !!

OpenStudy (amistre64):

using what we did for snafus post ....

OpenStudy (amistre64):

tangent line = point + n<tangent vector>

OpenStudy (amistre64):

and the plane is just the point stuck into the a plane equation using the normal vector

OpenStudy (anonymous):

wait i am still solving for T(t) and i get <-sin(t)/sec(t),cost/sect,-tant/sect> is it right so far ?

OpenStudy (amistre64):

1/sec = cos

OpenStudy (amistre64):

so yes so far :)

OpenStudy (anonymous):

ok thanks m still working on the furture steps !!

OpenStudy (anonymous):

now it asks to explain abt the motion of particle how would i know that ?

OpenStudy (amistre64):

a vector define a direction and a magnitude

OpenStudy (anonymous):

so (0,1,0)is the motion of particle ?

OpenStudy (amistre64):

in what direction is the particle moving? parallel to the Tangent vector, yes

OpenStudy (amistre64):

the Normal vector tells us in which direction the particle is turning as it moves